# A multidimensional measurement of photorealistic avatar quality of experience

### v2

@@ Abstract.
Photorealistic avatars are human avatars that look, move, and talk like real people. The performance of photorealistic avatars has significantly improved recently based on objective metrics such as PSNR, SSIM, LPIPS, FID, and FVD. However, recent photorealistic avatar publications do not provide subjective tests of the avatars to measure human usability factors. We provide an open source test framework to subjectively measure photorealistic avatar performance in ten dimensions: realism, trust, comfortableness using, comfortableness interacting with, appropriateness for work, creepiness, formality, affinity, resemblance to the person, and emotion accuracy. We show that the correlation of nine of these subjective metrics with PSNR, SSIM, LPIPS, FID, and FVD is weak, and moderate for emotion accuracy. The crowdsourced subjective test framework is highly reproducible and accurate when compared to a panel of experts. We analyze a wide range of avatars from photorealistic to cartoon-like and show that some photorealistic avatars are approaching real video performance based on these dimensions. We also find that for avatars above a certain level of realism, eight of these measured dimensions are strongly correlated. This means that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video. In addition, because there is a strong linear relationship between avatar affinity and realism, there is no uncanny valley effect for photorealistic avatars in the telecommunication scenario. We provide several extensions of this test framework for future work and discuss design implications for telecommunication systems. The test framework is available at https://github.com/microsoft/P.910.
@@ 1. Introduction
While there are many applications of photorealistic avatars, our focus is on evaluating them for the telecommunication scenario to improve trust and reduce video fatigue. Therefore the avatar test sequences we use are targeted for telecommunication, which includes people talking and expressing a wide range of emotions (happy, sad, surprised, fear, anger, and disgust).
@@ 2.1. Uncanny valley effect
The famous uncanny valley effect was first hypothesized by Mori in 1970 and later translated by MacDorman and Kageki (Mori et al., 2012), who described the phenomenon where humanoid objects that appear almost, but not quite, human evoke feelings of eeriness and discomfort in observers. Mori suggested when the affinity of humanoid objects is plotted on the y-axis against their human likeness on the x-axis, the affinity will increase as the likeness increases but then dips before going back up. This local dip in affinity is called the uncanny valley. For moving objects, Mori gave examples in order of human likeness: an industry robot, a humanoid robot, a zombie, a myoelectric hand, a bunraku puppet, an ill person, and a healthy person. The bottom of the uncanny valley occurred with the zombie and myoelectric hand examples. A meta-analysis of 72 experiments done on the uncanny valley effect is described in Diel et al (Diel et al., 2022), which used a three-level meta-analysis model and revealed the uncanny valley effect had a large effect size. Ho et al. (Ho and MacDorman, 2017) and Mathur et al. (Mathur and Reichling, 2016) were able to empirically reproduce the uncanny valley. Their studies included humanoid robots and the dip in affinity corresponded to images of robots. Tinwell (Tinwell et al., 2013) explores how avatar realism and facial expressions impact perceptions of eeriness and creepiness. Their results revealed that virtual characters that showed a lack of a startled response to a screaming sound were regarded as the most uncanny. Katsyri et al (Kätsyri et al., 2019) studied the uncanny valley using carefully matched images of virtual faces varying from artificial to realistic using both painted and computer-generated images. The painted images showed a linear relationship between affinity and human likeness, while the computer-generated showed more of an uncanny slope rather (flattening of the curve) than an uncanny valley.
@@ 2.2.1. Subjective evaluations of avatars
Canales et al. (Canales et al., 2024) conducted a study on the impact of avatar stylization and trust. Using the medical scenario with the avatar as the doctor, they measured trust, integrity, ability, and benevolence. While the style did not statistically significantly change these measurements, raters selected the mid-level avatar (between realistic and caricature) as their preferred doctor. However, this study only included the MetaHuman avatars and never measured the realism of these avatars. We found MetaHuman avatars have a realism MOS=2.3 (see Figure 6), which means the realistic avatars used in the study were not close to real video or SOTA photorealistic avatars. This may explain the different results reported by (Canales et al., 2024) and our present study, in which we show a strong relationship between realism and trust.
@@ 2.2.2. Subjective evaluation of video quality
A recent review of subjective image and video quality assessment tools is given by Testolina et al. (Testolina and Ebrahimi, 2021). ITU-T Rec. P.910 (P.910, 2021) provides a general subjective video quality assessment standard for multimedia applications. P.910 includes ACR (Absolute Category Rating), ACR-HR (ACR with Hidden Reference), DCR (Degradation Category Rating), and paired comparison (PC) methods, as well as rater qualifications, environment conditions, and video playback procedures. ITU-T Rec. P.911 is a counterpart of P.910 but for audiovisual signals. ITU-T Rec. P.912 (P.912, 2016) provides a target-specific subjective video quality assessment standard, such as for faces, license plates, etc. ITU-T Rec. P.913 (ITU-T Recommendation P.913, 2021) considers different displays and testing environments and provides flexibility on the rating scale and modality with mandatory reporting of test requirements (Pinson and Janowski, 2014). Recently, ITU-T Rec. P.911 and P.913 are merged into the P.910 providing a single recommendation given their large overlapping parts. ITU-T Rec. P.918 (ITU-T Recommendation P.918, 2021) details subjective assessment methods for five perceptual video quality dimensions, which can provide diagnostic information on the source of observed degradation. Finally, ITU-R BT.500 (ITU-R Recommendation BT.500-14, 2019) focuses on the video quality of broadcast television signals in a highly controlled environment.
@@ 2.2.3. Objective evaluations of avatars
Chen et al. (Chen et al., 2024) provide a dataset of photorealistic avatar videos from 36 subjects with the following transmission degradations: temporal artifacts, reduced texture resolution, and reduced frame rate. Each video has 26 ratings using ITU-T Rec. P.910 (P.910, 2021) to measure the avatar quality in the presence of these impairments. A new objective metric (HoloQA), is provided to predict subjective quality based on these impairments.
@@ 3.1. Measuring video quality
Measuring the quality of video is an important task in many engineering areas, such as video codec development, video enhancement, and video telecommunication systems. As a result, there are many standards and metrics for measuring video quality, though the gold standard is the subjective test done in a controlled lab environment. The most prevalent of these standardized subjective tests is the ITU-T P.910 (P.910, 2021). However, using P.910 in practice is slow due to the recruitment of test subjects and the limited number of test subjects, and expensive due to paying qualified test subjects and the cost of the test lab. The speed and cost result in the vast majority of research papers not using P.910 but rather objective functions that are not well correlated to subjective opinion.
@@ 3.1.1. Implementation
The implementation at https://github.com/microsoft/P.910 can be used either as an integrated survey in Amazon Mechanical Turk (AMT) or as an external survey deployed on a dedicated Web server. We also provide a lightweight container-based web application that can serve the experiment. The web application can easily be deployed on any Linux virtual machine. As a result, this implementation can be used for both crowdsourcing tests or remote testing with a dedicated panel of participants.
The open source implementation includes ACR, ACR-HR, DCR, and CCR. All methods can be used with either a five or nine-point discrete Likert scale. We also followed and extended best practices on video and speech quality assessment in crowdsourcing (Hossfeld et al., 2014; ITU-T PSTR-CROWDS, 2018) in our implementation.
@@ 3.1.2. Tools
We provide a set of program scripts to ease the interaction with the system and avoid operation errors. The scripts are used to create trapping sequences, process the submitted answers, and interact with AMT. Figure 1 shows the data flow diagram of the system.
Trapping sequences are customized to the dataset under the test and created by adding a text message asking participants to select a specific score. A test configuration and URLs for test sequences, trapping sequences, gold, and training clips are provided to the master script, which creates the HTML template of the test, a list of variables, and a configuration file for the result parser. Gold clips are video sequences whose quality is known to the experimenter. A test can be created in the HIT App Server by providing the HTML template and a list of variables. A generic project description and list of URLs can be downloaded and used to create a new test in AMT.
@@ 3.1.5. Qualification
Within the qualification section, the eligibility of test participants is evaluated. According to P.910, participants should be screened for normal color vision and normal or corrected-to-normal visual acuity (i.e., no error on the 20/30 line of a standard eye chart) (ITU-T Recommendation P.910, 2021). The standard Ishihara test for color vision (Clark, 1924) includes 15 plates in the normal and 6 plates in the short version, which are both too long for a crowdsourcing test. In a prestudy, we invited 300 participants from AMT (34% male) and 191 participants (91% male) from two internet communities dealing with color vision deficiencies. Both groups participated in the full Ishihara test in which 6% of participants from AMT and 96% from the online forums were detected as color blind. Applying a decision tree classifier with entropy as the criterion revealed that only using Plates 3 and 4 reached 98% accuracy (sensitivity 0.996, specificity 0.95). Consequently, we only use these two plates in the qualification section.
@@ 3.1.8. Training and rating
The training section includes videos from the training set, which cover the entire range of the rating scale. It also includes a trapping clip to test the participant’s understanding. Participants are alerted if they provide a wrong answer and asked to watch the video again. The training section is shown periodically to keep its anchoring effect, following best practices in the speech quality domain [(Naderi and Cutler, 2020)].
@@ 3.2. Avatar quality of experience survey design
The goal of the photorealistic avatar survey is to measure the quality of experience when using the avatar in multiple dimensions. We include realism, trust, and facial expressions (emotion accuracy) as suggested in ITU-T Rec. P.1320 (ITU-T Recommendation P.1320, 2022). We also include formality, creepiness, resemblance, appropriateness for work, comfortable using, and comfortable interacting with as used in (Inkpen and Sedlins, 2011). We added affinity to study the uncanny valley effect.
@@
Finally, in post processing, besides the answers to the above mentioned tests, outliers and variance in ratings will be checked.
@@ 4.2. Test set
We used 19 different avatars in the test set. The VASA-1 (Xu et al., 2024a), MS1, MS2, MS3, MS4 and MS5 avatars are from Microsoft. External1, External2, and External3 are 3 variations from the same organization. External1 is the AI model, External2 is the computer graphics based avatar and External3 is the set of avatars we created from the iPhone application from the organization. The Unreal [^1] https://www.unrealengine.com/en-US/metahuman avatars are models created using Epic’s Metahuman avatar creator. The Unreal_pro avatars are animations of Epic’s Metahuman store models created by professionals, while the non-pro Unreal avatars are created by the authors. The EMO (Tian et al., 2024), MegActor (Yang et al., 2024), Make-Your-Anchor (Huang et al., 2024), DisCo (Wang et al., 2024), Pose2Img (Qian et al., 2021), and Meta2 (Cao et al., 2022) avatars were scraped from their respective project pages. The Zoom avatar was a screen capture of the avatar in Zoom’s video conferencing application. The Meta1 avatar is from the Mark Zuckerberg, Lex Fridman interview [^2] https://www.youtube.com/watch?v=MVYrJJNdrEg. Audio2Photoreal and LivePortrait (Guo et al., 2024) were created using their open sourced models on GitHub. 9 of the 19 avatar models have more than one viewpoint variation (frontal, $45^{\circ}$ off-center and $90^{\circ}$ off-center). Thus we have a total of 40 different avatar conditions with an average of 5.4 clips per variation for Template A. This also includes the baseline clips of real recordings of the actors. For Template B, we only used models that had input video clips available, resulting in 26 different avatar conditions from 13 different avatars with an average of 5 clips per variation. Table 5 provides the number of people for each avatar, the number viewpoints, whether speech is included, and if there was a reference video driving the avatar.

### v3

@@ Abstract.
Photorealistic avatars are human avatars that look, move, and talk like real people. The performance of photorealistic avatars has significantly improved recently based on objective metrics such as PSNR, SSIM, LPIPS, FID, and FVD. However, recent photorealistic avatar publications do not provide subjective tests of the avatars to measure human usability factors. We provide an open source test framework to subjectively measure photorealistic avatar performance in ten dimensions: realism, trust, comfortableness using, comfortableness interacting with, appropriateness for work, creepiness, formality, affinity, resemblance to the person, and emotion accuracy. Using telecommunication scenarios, we show that the correlation of nine of these subjective metrics with PSNR, SSIM, LPIPS, FID, and FVD is weak, and moderate for emotion accuracy. The crowdsourced subjective test framework is highly reproducible and accurate when compared to a panel of experts. We analyze a wide range of avatars from photorealistic to cartoon-like and show that some photorealistic avatars are approaching real video performance based on these dimensions. We also find that for avatars above a certain level of realism, eight of these measured dimensions are strongly correlated. This means that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video. In addition, because there is a strong linear relationship between avatar affinity and realism, there is no uncanny valley effect for photorealistic avatars in the telecommunication scenario. We suggest several extensions of this test framework for future work and discuss design implications for telecommunication systems. The test framework is available at https://github.com/microsoft/P.910.
@@ 1. Introduction
While there are many applications of photorealistic avatars, our focus is on evaluating them for the telecommunication scenario to improve trust and reduce video fatigue. Therefore the avatar test sequences we use are targeted for telecommunication, which includes people talking and expressing a wide range of emotions (happy, sad, surprised, fear, anger, and disgust). However, we think that the test framework and methodology is applicable to all of the above mentioned avatar applications.
@@ 2.1. Uncanny valley effect
The famous uncanny valley effect was first hypothesized by Mori in 1970 and later translated by MacDorman and Kageki (Mori et al., 2012), who described the phenomenon where humanoid objects that appear almost, but not quite, human evoke feelings of eeriness and discomfort in observers. Mori suggested when the affinity of humanoid objects is plotted on the y-axis against their human likeness on the x-axis, the affinity will increase as the likeness increases but then dips before going back up. This local dip in affinity is called the uncanny valley. For moving objects, Mori gave examples in order of human likeness: an industry robot, a humanoid robot, a zombie, a myoelectric hand, a bunraku puppet, an ill person, and a healthy person. The bottom of the uncanny valley occurred with the zombie and myoelectric hand examples. A meta-analysis of 72 experiments done on the uncanny valley effect is described in Diel et al. (Diel et al., 2022), which used a three-level meta-analysis model and revealed the uncanny valley effect had a large effect size. Ho et al. (Ho and MacDorman, 2017) and Mathur et al. (Mathur and Reichling, 2016) were able to empirically reproduce the uncanny valley. Their studies included humanoid robots and the dip in affinity corresponded to images of robots. Tinwell (Tinwell et al., 2013) explores how avatar realism and facial expressions impact perceptions of eeriness and creepiness. Their results revealed that virtual characters that showed a lack of a startled response to a screaming sound were regarded as the most uncanny. Katsyri et al (Kätsyri et al., 2019) studied the uncanny valley using carefully matched images of virtual faces varying from artificial to realistic using both painted and computer-generated images. The painted images showed a linear relationship between affinity and human likeness, while the computer-generated showed more of an uncanny slope rather (flattening of the curve) than an uncanny valley.
@@ 2.2.1. Subjective evaluations of avatars
Canales et al. (Canales et al., 2024) conducted a study on the impact of avatar stylization and trust. Using the medical scenario with the avatar as the doctor, they measured trust, integrity, ability, and benevolence. While the style did not statistically significantly change these measurements, raters selected the mid-level avatar (between realistic and caricature) as their preferred doctor. However, this study only included the MetaHuman avatars and never measured the realism of these avatars. We found MetaHuman avatars have a realism mean opinion score (MOS)=2.3 (see Figure 6), which means the realistic avatars used in the study were not close to real video or SOTA photorealistic avatars. This may explain the different results reported by (Canales et al., 2024) and our present study, in which we show a strong relationship between realism and trust.
@@
Gasch et al. (Gasch et al., 2024) examined how using avatars of varying realism in virtual reality video conferencing affects users’ sense of presence, privacy concerns, and acceptance of the technology for work and education. They find that hyper-realistic avatars generate a stronger sense of presence and higher user preference, although privacy concerns increase notably if another person uses someone else’s avatar.
Frampton-Clerk et al. (Frampton-Clerk and Oyekoya, 2022) explored how people perceive the realism of someone else’s look-alike avatar, focusing on key features like lip syncing, facial expression, and full body movement. They show that while full facial and body animations enhance realism, having only lip syncing can trigger unsettling (uncanny valley) reactions.
Garau et al. (Garau et al., 2003) investigated how avatar realism and eye gaze control affect perceived communication quality in a shared immersive virtual environment. They show that matching higher avatar realism with a more realistic gaze model significantly enhances users’ sense of presence and face-to-face interaction, whereas lower-realism avatars do not benefit from advanced gaze behaviors.
Suk et al. (Suk and Laine, 2023) studied how avatar facial appearance in immersive virtual reality (VR) affects users’ perceived embodiment and presence. They find that avatars with facial features resembling the user’s own face significantly enhance the sense of embodiment but have no substantial effect on the user’s sense of presence.
Junuzovic et al. (Junuzovic et al., 2012) compared avatar, video, and audio conferencing for four-person workplace meetings and found that video conferencing was rated highest across all dimensions, including social presence, realism, and usefulness, due to superior non-verbal cues. Although avatars provided more sociability and a shared virtual space compared to audio alone, participants considered cartoon avatars inappropriate for professional settings and expressed frustrations with tracking inaccuracies, but noted avatar conferencing had potential if perfected.
Lin et al. (Lin et al., 2023) explores the effectiveness of various visual indicators (such as color coding, scanning effects, transparency, and badges) to convey the authenticity of photorealistic avatars in social VR, and their impacts on perceived trustworthiness. They conclude that explicitly displaying the avatar user’s name enhances perceived trust, while most other visual indicators negatively affect trust, particularly as interactivity and immersion levels increase.
Wang et al. (Wang et al., 2024b) explores using life-size 2D video-based avatars in augmented reality (AR) teleconferencing on head-mounted displays (HMDs) to balance visual realism and the sense of co-presence. Through pilot studies and user evaluations, the authors identify optimal avatar placements for small-group conversations and demonstrate that 2D video avatars provide a superior balance between visual fidelity and co-presence compared to traditional methods, also highlighting how the Field of View (FoV) affects optimal avatar placement.
Yu et al. (Yu et al., 2021) compares two avatar embodiment techniques—point cloud reconstruction (PCR) avatars and virtual character-based avatars (3DVC)—in an asymmetric VR/AR teleconsultation system. They showed that despite occlusions and missing facial details, PCR avatars provided superior perceptions of copresence, social presence, behavioral realism, and humanness compared to virtual character avatars, although neither method significantly affected objective task performance.
@@ 2.2.2. Subjective evaluation of video quality
A recent review of subjective image and video quality assessment tools is given by Testolina et al. (Testolina and Ebrahimi, 2021). ITU-T Rec. P.910 ({ITU-T Recommendation P.910}, 2021) provides a general subjective video quality assessment standard for multimedia applications. P.910 includes ACR (Absolute Category Rating), ACR-HR (ACR with Hidden Reference), DCR (Degradation Category Rating), and paired comparison (PC) methods, as well as rater qualifications, environment conditions, and video playback procedures. ITU-T Rec. P.911 is a counterpart of P.910 but for audiovisual signals. ITU-T Rec. P.912 (P.912, 2016) provides a target-specific subjective video quality assessment standard, such as for faces, license plates, etc. ITU-T Rec. P.913 (ITU-T Recommendation P.913, 2021) considers different displays and testing environments and provides flexibility on the rating scale and modality with mandatory reporting of test requirements (Pinson and Janowski, 2014). Recently, ITU-T Rec. P.911 and P.913 are merged into the P.910 providing a single recommendation given their large overlapping parts. ITU-T Rec. P.918 (ITU-T Recommendation P.918, 2021) details subjective assessment methods for five perceptual video quality dimensions, which can provide diagnostic information on the source of observed degradation. Finally, ITU-R BT.500 (ITU-R Recommendation BT.500-14, 2019) focuses on the video quality of broadcast television signals in a highly controlled environment.
@@ 2.2.3. Objective evaluations of avatars
Chen et al. (Chen et al., 2024) provide a dataset of photorealistic avatar videos from 36 subjects with the following transmission degradations: temporal artifacts, reduced texture resolution, and reduced frame rate. Each video has 26 ratings using ITU-T Rec. P.910 ({ITU-T Recommendation P.910}, 2021) to measure the avatar quality in the presence of these impairments. A new objective metric (HoloQA), is provided to predict subjective quality based on these impairments.
@@ 2.3. Trust, empathy, and fatigue for telecommunication
2.4. Virtual reality and kinetic embodiment
Biehl et al. (Biehl et al., 2015) showed that local participants exhibited significantly more overlapping talk with remote participants who used an embodied proxy, than with remote participants in basic-video conferencing. In addition, while kinetic embodied technology increased the local participants’ perceived presence of remote teammates, it did not enhance remote participants’ own sense of telepresence.
Ratan et al.(Ratan and Hasler, 2014) explores how people present themselves and perceive others in social VR, highlighting a strong preference among users to keep avatars consistent with their physical selves and a heightened importance of voice in verifying identity cues (such as gender, age, and ethnicity). It further shows that social VR can influence how users understand their own identities—especially for those exploring or affirming marginalized or changing self-images—through the immersive and embodied nature of virtual interactions.
Freeman et al. (Freeman and Maloney, 2021) showed people interact socially in VR using many of the same verbal and nonverbal cues familiar from face-to-face (F2F) settings, but they deploy them differently: VR users rely more on direct verbal strategies (like immediately introducing a topic) and simpler gestures (like waving) than they do in F2F encounters. In contrast, F2F interactions more frequently involve subtle cues like moving closer or making small sounds to get someone’s attention, indicating that VR changes both the norms and practices of social engagement.
Toothman et al. (Toothman and Neff, 2019) investigated how different types of avatar tracking errors in Virtual Reality—such as latency, vibration, popping, and stuttering—affect user experiences including performance, embodiment, enjoyment, usability, and spatial presence. While substantial tracking errors negatively impacted task performance and users’ sense of embodiment, they did not significantly reduce users’ social presence during virtual interactions, even at extreme error levels.
@@ 3.1. Measuring video quality
Measuring the quality of video is an important task in many engineering areas, such as video codec development, video enhancement, and video telecommunication systems. As a result, there are many standards and metrics for measuring video quality, though the gold standard is the subjective test done in a controlled lab environment. The most prevalent of these standardized subjective tests is the ITU-T P.910 ({ITU-T Recommendation P.910}, 2021). However, using P.910 in practice is slow due to the recruitment of test subjects and the limited number of test subjects, and expensive due to paying qualified test subjects and the cost of the test lab. The speed and cost result in the vast majority of research papers not using P.910 but rather objective functions that are not well correlated to subjective opinion.
@@ 3.1.1. Implementation
The implementation at https://github.com/microsoft/P.910 can be used either as an integrated survey in Amazon Mechanical Turk (MTurk) or as an external survey deployed on a dedicated Web server. We also provide a lightweight container-based web application that can serve the experiment. The web application can easily be deployed on any Linux virtual machine. As a result, this implementation can be used for both crowdsourcing tests or remote testing with a dedicated panel of participants.
The open source implementation includes ACR, ACR-HR, DCR, and CCR. All methods can be used with either a five or nine-point discrete Likert scale. We also followed and extended best practices on video and speech quality assessment in crowdsourcing (Hossfeld et al., 2014; ITU-T PSTR-CROWDS, 2018) in our implementation. The avatar quality of experience uses ACR and CCR measurements, and extensions could use ACR-HR and DCR.
@@ 3.1.2. Tools
We provide a set of program scripts to ease the interaction with the system and avoid operation errors. The scripts are used to create trapping sequences, process the submitted answers, and interact with MTurk. Figure 1 shows the data flow diagram of the system.
Trapping sequences are customized to the dataset under the test and created by adding a text message asking participants to select a specific score. A test configuration and URLs for test sequences, trapping sequences, gold, and training clips are provided to the master script, which creates the HTML template of the test, a list of variables, and a configuration file for the result parser. Gold clips are video sequences whose quality is known to the experimenter. A test can be created in the HIT App Server by providing the HTML template and a list of variables. A generic project description and list of URLs can be downloaded and used to create a new test in MTurk.
@@ 3.1.5. Qualification
Within the qualification section, the eligibility of test participants is evaluated. According to P.910, participants should be screened for normal color vision and normal or corrected-to-normal visual acuity (i.e., no error on the 20/30 line of a standard eye chart) (ITU-T Recommendation P.910, 2021). The standard Ishihara test for color vision (Clark, 1924) includes 15 plates in the normal and 6 plates in the short version, which are both too long for a crowdsourcing test. In a prestudy, we invited 300 participants from MTurk (34% male) and 191 participants (91% male) from two internet communities dealing with color vision deficiencies. Both groups participated in the full Ishihara test in which 6% of participants from MTurk and 96% from the online forums were detected as color blind. Applying a decision tree classifier with entropy as the criterion revealed that only using Plates 3 and 4 reached 98% accuracy (sensitivity 0.996, specificity 0.95). Consequently, we only use these two plates in the qualification section.
@@ 3.1.8. Training and rating
The training section includes videos from the training set, which cover the entire range of the rating scale. It also includes a trapping clip to test the participant’s understanding. Participants are alerted if they provide a wrong answer and asked to watch the video again. The training section is shown periodically to keep its anchoring effect, following best practices in the speech quality domain (Naderi and Cutler, 2020).
@@ 3.2. Avatar quality of experience survey design
The goal of the photorealistic avatar survey is to measure the quality of experience when using the avatar in multiple dimensions. We include realism, trust, and facial expressions (emotion accuracy) as suggested in ITU-T Rec. P.1320 (ITU-T Recommendation P.1320, 2022) to measure the quality of experience. We also include formality, creepiness, resemblance, appropriateness for work, comfortable using, and comfortable interacting with as used in (Inkpen and Sedlins, 2011). We added affinity to study the uncanny valley effect, which is the simplest dimension that has been used to validate the uncanny valley (Mathur and Reichling, 2016) and is what is used in the original hypothesis of the uncanny valley (Mori et al., 2012).
@@
Finally, in post processing, besides the answers to the above mentioned tests, outliers and variance in ratings will be checked. The accepted submissions are aggregated per clip and per model, and MOS, standard deviation, and 95% confidence interval for each item will be reported.
@@ 4.2. Test set
We used 19 different avatars in the test set. The VASA-1 (Xu et al., 2024a), MS1, MS2, MS3, MS4 and MS5 avatars are from Microsoft. External1, External2, and External3 are 3 variations from the same organization. External1 is the AI model, External2 is the computer graphics based avatar and External3 is the set of avatars we created from the iPhone application from the organization. The Unreal [^1] https://www.unrealengine.com/en-US/metahuman avatars are models created using Epic’s Metahuman avatar creator. The Unreal_pro avatars are animations of Epic’s Metahuman store models created by professionals, while the non-pro Unreal avatars are created by the authors. The EMO (Tian et al., 2024), MegActor (Yang et al., 2024), Make-Your-Anchor (Huang et al., 2024), DisCo (Wang et al., 2024a), Pose2Img (Qian et al., 2021), and Meta2 (Cao et al., 2022) avatars were scraped from their respective project pages. The Zoom avatar was a screen capture of the avatar in Zoom’s video conferencing application. The Meta1 avatar is from the Mark Zuckerberg, Lex Fridman interview [^2] https://www.youtube.com/watch?v=MVYrJJNdrEg. Audio2Photoreal and LivePortrait (Guo et al., 2024) were created using their open sourced models on GitHub. 9 of the 19 avatar models have more than one viewpoint variation (frontal, $45^{\circ}$ off-center and $90^{\circ}$ off-center). Thus we have a total of 40 different avatar conditions with an average of 5.4 clips per variation for Template A. This also includes the baseline clips of real recordings of the actors. For Template B, we only used models that had input video clips available, resulting in 26 different avatar conditions from 13 different avatars with an average of 5 clips per variation. Table 5 provides the number of people for each avatar, the number viewpoints, whether speech is included, and if there was a reference video driving the avatar.
This test set should help answer RQ1, as we can compute objective and subjective metrics on it and compute their correlation. We can also use this test set to answer RQ2, as we can determine the importance and redundancy of each dimension measured in the survey. RQ3 can be answered by using this test set to show the results are both accurate compared to experts and reproducible. By analyzing affinity and realism from this test set we can determine if we see a dip in affinity as realism increases, which answers RQ4. Finally, because we include real videos in this test set (which we consider as the baseline), we can determine the gaps between real video and avatars in each dimension.